OpenStudy (anonymous):

The speed of sound in the air is proportional to the square root of the absolute temperature. At 20 °C the speed of sound is 343 m/s. What is the speed of sound at –10 °C? (0 °C = 273 K).

OpenStudy (nowhereman):

Just solve: \[ \frac{x}{\sqrt{273 - 10}} = \frac{343}{\sqrt{273+20}} \] because that is what proportional means.

OpenStudy (anonymous):

how do you get kelvin from celcius?

OpenStudy (nowhereman):

It's just a simple translation [K] = [C] + 273.15, but you should see this from my previous answer.

OpenStudy (anonymous):

am i not supposed to use something like \[(spped of sound) \alpha temperature ^{2}\] ?

OpenStudy (anonymous):

speed = k * temperature ^ 2

OpenStudy (anonymous):

?

OpenStudy (nowhereman):

No, you said yourself, that the speed of sound v is proportional to the square root of the temperature T, so \[v \sim \sqrt{T} ⇔ ∃ k: v = k\cdot \sqrt{T}\]

OpenStudy (anonymous):

oh, i get it

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

so, is the answer 325? because thats what i got

OpenStudy (anonymous):

anwar, can you please fan lokisan, because he helped me alot and i can only fan him once

OpenStudy (nowhereman):

Yes, 325 is about right. A more exact value I got, using 1 deg C = 273.15 K, is 324.975677595.